### v3

# Calibration in Deep Learning: A Survey of the State-of-the-Art

Abstract
Calibrating deep neural models plays an important role in building reliable, robust AI systems in safety-critical applications. Recent work has shown that modern neural networks that possess high predictive capability are poorly calibrated and produce unreliable model predictions. Though deep learning models achieve remarkable performance on various benchmarks, the study of model calibration and reliability is relatively under-explored. Ideal deep models should have not only high predictive performance but also be well calibrated. There have been some recent advances in calibrating deep models. In this survey, we review the state-of-the-art calibration methods and their principles for performing model calibration. First, we start with the definition of model calibration and explain the root causes of model miscalibration. Then we introduce the key metrics that can measure this aspect. It is followed by a summary of calibration methods that we roughly classify into four categories: post-hoc calibration, regularization methods, uncertainty estimation, and composition methods. We also cover recent advancements in calibrating large models, particularly large language models (LLMs). Finally, we discuss some open issues, challenges, and potential directions.
1 Introduction
Deep Neural Networks (DNNs) have been showing promising predictive power in many domains such as computer vision (?), speech recognition (?) and natural language processing (?). Nowadays, deep neural network models are frequently being deployed into real-world systems. However, recent work (?) pointed out that those highly accurate, negative-log-likelihood (NLL) trained deep neural networks are poorly calibrated (?), i.e., the model predicted class probabilities do not faithfully estimate the true correctness likelihood and lead to overconfident and underconfident predictions.
Although in production environments, the primary goal of deep learning models is to make accurate predictions that can inform decision-making processes, in many real-world applications, decisions are not based solely on whether a prediction is correct or incorrect, but also on the confidence associated with that prediction. Deploying uncalibrated models into real-world systems is at high risk, particularly for safety-critical applications such as medical diagnosis (?), autonomous driving (?) and finance decision-making (?). One concrete example is the decision making based on model predictions in a medical diagnostic system. The model might predict that a patient has a 70% chance of having a certain disease based on their symptoms and medical history. A well-calibrated model ensures that if it predicts a 70% chance of an event happening, that event indeed occurs approximately 70% of the time in reality. Without calibration, the reliability of model predictions is compromised, leading to inaccurate assessments of uncertainty and potentially flawed decision-making.
Figure 1: An illustration of model calibration. Uncalibrated model (left) that trained with standard cross-entropy loss and calibrated model (right) that trained with focal loss ($\gamma=5$), have similar predictive performance on a binary classification task (accuracy is 83.8% and 83.4% respectively), but the right one is better calibrated. Top: The reliability diagram plots with 10 bins. The diagonal dash line presents perfect calibration (on a specific bin, confidence is equal to accuracy.), Expected Calibration Error (ECE) and Maximum Calibration Error (MCE) are used to measure model calibration performance. Bottom: the posterior distribution of the two models on 1000 samples, the one from calibrated model is better distributed.
Calibrating deep models is a procedure for preventing the model’s posterior distribution from being over- or under-confident. Figure 1 gives an illustration of calibrating a binary classification model. It is noted that (1) a highly predictive model can be poorly calibrated, this is exhibited by high calibration errors; (2) The deep models tend to be primarily overconfident, this is shown by a spiking posterior distribution  (?, ?, ?). Model miscalibration is caused by many factors such as over-parameterized networks, a lack of appropriate regularization, limited data, imbalanced label distributions, etc. We discuss those factors in details in the Section 2.2.
1.1 Scope and Focus
In recent years, a variety of approaches have emerged for model calibration, spanning post-hoc adjustments, regularization techniques, uncertainty estimation, and hybrid methods. Several related surveys have delved into this area (?), or those focusing on the highly relevant topic of uncertainty estimation (?, ?, ?), which briefly touch upon model calibration. However, there remains a gap in the literature regarding a comprehensive review of recently proposed calibration techniques. Setting itself apart from previous surveys, our research emphasizes several key distinctions:
  • This survey reviews the state-of-the-art calibration methods and focuses mostly on the ones proposed in recent years. Those methods such as kernel-based methods, differentiable calibration proxy, and meta-learning-based approaches. Those are rarely discussed in previous and existing surveys.
  • This survey tries to explain calibration principles of each method via the discussion of the conceptual relationships among over-parameterization, over-fitting, and over-confidence. We systematically categorize those methods into post-hoc, regularization (explicit, implicit and differentiable calibration proxy), uncertainty estimation, and hybrid methods that combines multiple calibration techniques.
  • This survey also delves into the methodologies for calibrating large pre-trained models, with a specific focus on large language models (LLMs). Notably, the calibration of LLMs for zero-shot inference has garnered escalating interest within AI communities.
The rest of this survey is structured as follows.
  • Section 2 provides an introduction to the concept of model calibration, elucidating the factors that contribute to miscalibration. We explore how various aspects such as model complexity, data distribution, training and measurement procedures can impact the calibration of machine learning models.
  • Section 3 presents an overview of mainstream calibration metrics commonly used to evaluate the calibration performance of models. We discuss the significance of metrics such as expected calibration error (ECE), reliability diagrams, and other calibration metrics in assessing the alignment between predicted probabilities and true outcomes.
  • Section 4 delves into a comprehensive review and classification of recently proposed calibration methods. We categorize these methods based on their underlying principles and techniques, including post-hoc adjustments, regularization approaches, uncertainty estimation techniques, and novel hybrid methods. For each category, we provide in-depth discussions, highlighting their strengths and limitations.
  • Section 5 discusses potential future directions in the field of model calibration. We identify emerging trends, research gaps, and challenges that warrant further investigation. Additionally, we offer insights into promising avenues for advancing calibration techniques.
2 Preliminaries and Backgrounds
This section describes the definition of model calibration and the aspects cause miscalibration.
2.1 Definitions
In classification tasks, for a given input variable $X$ and a categorical variable $Y\in\{1,2,...,k\}$, assume we have a neural network model $f$ which maps input variable $\mathbf{x}$ to a categorical distribution $p=\{p_{1},...,p_{k}\}$ over $k$ classes $\{y_{1},...,y_{k}\}$: $f:D\rightarrow\Delta$, where $\Delta$ is the $k-1$ dimensional standard probability simplex and $\Delta=\{p\in[0,1]^{k}|\sum_{i=1}^{k}p_{i}=1\}$. Calibration measures the degree of the match between predicted probability $p$ and the true correctness likelihood. A model $f$ is perfectly calibrated on if and only if:
$\mathbb{P}(Y=y_{i}|f(X)=p)=p_{i}$ (1)
where $p$ is true correctness likelihood. Intuitively, for all input pairs $\{x,y\}\in D$, if model predicts $p_{i}=0.8$, we expect that 80% have $y_{i}$ as label.
Instead of probability distribution, the argmax calibration (?, ?, ?) takes only the maximum probability into consideration:
$\mathbb{P}(Y\in\arg\max(p)|\max(f(X))=p^{*})=p^{*}$ (2)
In reality, it is difficult to obtain perfect calibration, any deviation from it represents miscalibration.
2.2 Aspects Impact Model Calibration
It has been observed that some recent changes in modern neural networks are responsible for model miscalibration (?, ?, ?). The underlying general cause is that modern neural networks’ high capacity makes them vulnerable to miscalibration, which is tightly correlated to the concepts of over-parameter, overfitting and over-confidence. Figure 2 outlines those factors that impact model calibration.
Figure 2: Illustration depicting factors influencing model calibration. Modern neural networks often exhibit overconfidence, which is closely associated with overfitting during model training. Overfitting is commonly exacerbated by inadequate regularization techniques and the presence of data bias within training datasets, particularly in over-parameterized models.
2.2.1 Model Complexity
While increasing the depth and width of neural networks helps to obtain highly predictive power, it also negatively increases calibration errors. Empirical evidence has shown that this poor calibration is linked to overfitting on the negative log-likelihood (NLL) (?, ?). The over-parameteriaztion is one of the main causes of overfitting. Concretely, along with the standard NLL-based model training, when classification error is minimized, keeping training will further push the model to minimize NLL on the training data, i.e., push the predicted softmax probability distribution as close as possible to the ground-truth distribution (which is usually one-hot representation). Model overfitting starts by exhibiting increased test NLL, and then the model becomes overconfident (?). This phenomena becomes more severe when training a high capacity model with limited model.
2.2.2 Data Issues
Another important aspects that directly impact model calibration are: data quantity, i.e., the scale of dataset and data quality such as relevance, consistency, completeness, etc. Training high-capacity (over-parameterized) networks with scarce data can easily cause overfitting and have an overconfident model. Data augmentation is an effective way to alleviate this phenomenon and brings implicit calibration effects (?, ?). The recent pretrain-finetun paradigm offers the possibility of reducing overfitting caused by limited data (?). Besides having a decent amount of training data, we should also highlight the importance of addressing the following issues:
  • Data noise: this refers to the presence of erroneous or misleading information in datasets, which can adversely affect model training and performance. Noisy data can arise from various sources such as systematic or measurement errors when pulling historical data from production models, labeling or annotation mistakes caused by human annotators, or outliers  (?). Those factors can lead to biased model estimates, reduced predictive accuracy, and poor generalization to unseen data. This contributes to model miscalibration. Some methods has been proposed to mitigate the impact of data noise on model calibration such as robust loss functions or adversarial training, have shown promise in improving model robustness to noisy data (?, ?). Also Ensemble methods have been proofed effective in identifying and mitigating the effects of noisy data on model performance (?, ?).
  • Data imbalance. It can significantly impact model calibration. When a dataset is imbalanced, meaning that one class or outcome is significantly more prevalent than others, it can lead to biased predictions and poor calibration. This is because the model may become overly confident in predicting the majority class while being less certain about the minority classes. Focal loss  (?, ?) has shown its effectiveness in calibrating models that trained with imbalanced datasets (?).
  • Data biases and fairness. Extensive studies have highlighted the presence of biases in datasets. Biases arises due to historical disparities, societal prejudices, or systemic inequalities. For instance, facial recognition datasets often exhibit racial and gender biases (?). Mitigating bias and ensuring fairness are imperative steps towards building models that exhibit better calibration and reliability (?, ?). This involves a multifaceted approach that begins with critically examining the data collection process to identify and rectify sources of bias. Techniques such as data augmentation (?, ?), bias mitigation algorithms (?), and fairness-aware model training (?) can help mitigate biases during the model development stage.
2.2.3 Learning Objective and Regularization
The choice of loss function and optimization objective can significantly impact the calibration quality of machine learning models. For instance, cross-entropy loss is commonly used for classification tasks, as it directly optimizes for the likelihood of correct class predictions. However, optimizing solely for accuracy may lead to poorly calibrated probability estimates, especially in situations where class distributions are imbalanced or ambiguous. Focal loss (?, ?), as alternative loss, has recently demonstrated promising performance in calibrating deep models.
Regularization can effectively prevent overfitting when model capacity increases. Recent trends suggest that explicit L2 regularization may not be necessary to achieve a highly accurate model when applying batch normalization (?) or dropout (?), but Guo et al. (?) empirically demonstrated model tends to be less calibrated without using L2 regularization. There have been more recent regularization techniques  (?, ?, ?, ?) been proposed to improve model calibration. We will discuss those methods in more details in the Section 4.
2.2.4 Bias in Evaluation Metrics
Another important factor that does not directly cause model calibration, but hurts accurate assessment of model calibration is the introduced bias in evaluation metrics. For instance, the bias that introduced by binning mechanisms in the widely used Expected Calibration Error (ECE) has been identified in  (?). By binning continuous variables, fine-grained details in the data may be lost, leading to an oversimplified representation of the underlying relationships. This loss of information can result in calibration bias, as the model’s predicted probabilities may not accurately capture the nuances of the data distribution.
To mitigate calibration bias introduced by binning mechanisms , it is important to carefully consider the choice of bin boundaries and the impact of discretization on the model’s calibration performance. There have been some techniques proposed to address this issues such as Adaptive ECE (?), Smooth ECE (?). We will further discuss in the Section 4.
3 Calibration Measurements
Exact calibration measurement with finite data samples is impossible given that the confidence $p$ is a continuous variable (?). There are some popular metrics that approximate model calibration error by grouping $N$ predictions into $M$ interval bins $\{b_{1},b_{2},...,b_{M}\}$.
3.0.1 Expected Calibration Error (ECE)
ECE (?) is a scalar summary statistic of calibration. It quantifies the discrepancy between predicted probabilities and observed frequencies by dividing the prediction space into several bins based on the predicted probabilities. Then, within each bin, it calculates the average absolute difference between the predicted probabilities and the empirical frequencies of the events occurring. The ECE is computed as the weighted average of these differences across all bins. Mathematically, ECE can be expressed as:
$\textsc{ECE}=\frac{1}{N}\sum_{m=1}^{M}\left|b_{m}\right||\textrm{acc}(b_{m})-\textrm{conf}(b_{m})|$ (3)
where $N$ is the total number of samples. $\left|b_{m}\right|$ is the number of samples in bin $b_{m}$, and
$\textrm{acc}(b_{m})=\frac{1}{\left|b_{m}\right|}\sum_{i\in B_{m}}\mathbf{1}(\hat{y}_{i}=y_{i}),~{}\textrm{conf}(b_{m})=\frac{1}{\left|b_{m}\right|}\sum_{i\in B_{m}}p_{i}.$ (4)
3.0.2 Maximum Calibration Error (MCE)
MCE (?), MCE focuses on identifying the largest discrepancy between predicted probabilities and empirical accuracies within any individual bin,
$\textsc{MCE}=\max_{m\in\{1,\dots,M\}}|\textrm{acc}(b_{m})-\textrm{conf}(b_{m})|.$ (5)
and is particularly important in high-risk applications where reliable confidence measures are absolutely necessary. While MCE highlights the most substantial calibration error observed within the prediction space, it may not offer as comprehensive an assessment of overall calibration performance as ECE does. ECE considers the average discrepancy across all bins, providing a more holistic evaluation of calibration quality.
3.0.3 Classwise ECE (CECE)
Classwise ECE (?) can be seen as the macro-averaged ECE. It extends the bin-based ECE to measure calibration across all the possible $K$ classes. In practice, predictions are binned separately for each class, and the calibration error is computed at the level of individual class-bins and then averaged. The metric can be formulated as
$\textsc{CECE}=\sum_{m=1}^{M}\sum_{c=1}^{K}\frac{|b_{m,c}|}{NK}|\textrm{acc}_{c}(b_{m,c})-\textrm{conf}_{c}(b_{m,c})|$ (6)
where $b_{m,c}$ represents a single bin for class $c$. In this formulation, $\textrm{acc}_{c}(b_{m,c})$ represents average binary accuracy for class $c$ over bin $b_{m,c}$ and $\textrm{conf}_{c}(b_{m,c})$ represents average confidence for class $c$ over bin $b_{m,c}$.
3.0.4 Adaptive ECE (AECE)
The binning mechanism in the aforementioned metrics can introduce bias; the pre-defined bin size determines the number of samples in each bin. Adaptive ECE (?) introduces a new binning strategy to use an adaptive scheme that spaces the bin intervals to ensure each bin hosts an equal number of samples.
$\textsc{AECE}=\sum_{r=1}^{R}\sum_{c=1}^{K}\frac{1}{RK}|\textrm{acc}_{c}(b_{n,c})-\textrm{conf}_{c}(b_{n,c})|$ (7)
Where $r\in[1,R]$ is defined by the $[N/R]$-th index of the sorted and threshold predictions.
For a perfectly calibrated classifier, those calibration erros should equal 00.
3.0.5 Reliability Diagram
Besides the metrics that provide a scalar summary on calibration, reliability diagram (also referred to as a calibration plot or calibration curve 1) visualizes whether a model is over- or under-confident on bins by grouping predictions into bins according to their prediction probability. This visual representation enables a detailed examination of how well the predicted probabilities align with the observed frequencies of outcomes, providing insights into the reliability of the model’s confidence estimates.The diagonal line in Figure 1 presents perfect calibration:
$\textrm{acc}(b_{m})=\textrm{conf}(b_{m}),\forall m$ (8)
The red bar presents the gap to perfect calibration. Deviations from the perfect calibration line indicate calibration errors. Points above the diagonal suggest that the model’s predicted probabilities are overly confident, while points below the diagonal indicate underconfidence. The magnitude and direction of these deviations provide valuable insights into the systematic biases present in the model’s probabilistic predictions.
4 Calibration Methods
In this section, we categorize the state-of-the-art calibration methods into post-hoc methods, regularization methods and uncertainty estimation methods. Besides, we discuss hybrid methods that combine different calibration methods. Figure 3 summarizes and categories those methods.
{forest}
for tree=draw, semithick, rounded corners, edge = draw, semithick, anchor = east, grow = east, forked edge, s sep = 2mm, l sep = 2mm, fork sep = 2mm, [Calibration [Large Language Models [Contextual Calibration (?)] [Prototypical Calibration (?)] [Domain-Context Calibration (?)] [Batch Calibration (?)] ] [Hybird Methods [LS + Mixup (?)] [Ensemble TS (ETS) (?)] [TS + MC dropout (?)] ] [ Uncertainty Estimation [Bayesian Methods [Finite-State Probabilistic RNNs (?)] [MC Dropout (?)] [Bayesian RNNs (?)] [BNNs (?)] ] [Ensemble Methods [Deep Ensembles (?)] ] ] [Regularization Methods [Explicit Regularization [Huber loss (?)] [DCA (?)] [Entropy Regularization (?)] ] [Implicit Regularization [Focal Loss (?)] [FLSD (?)] [Label Smooth(LS) (?)] [Mixup (?)] ] [Differentiable Proxy [MMCE (?)] [Meta-Calibration (?)] [AvUC (?)] [MDCA (?)] ] ] [Post-hoc Methods [Local TS (LTS) (?)] [Dirichlet calibration (?)] [Bin-wise TS (BTS) (?)] [Attended TS (ATS) (?)] [Temperature Scaling (TS) (?)] ] ]
Figure 3: The categorization of calibration methods and the representative methods in each category. The regularization methods are further divided into explicit, implicit regularization and differentiable calibration proxies.
4.1 Post-hoc Methods
Post-hoc calibration methods aim to calibrate a model after training. Those include non-parametric calibration histogram binning(?), isotonic regression (?) and parametric methods such as Bayesian binning into quantiles (BBQ) and Platt scaling (?). Out of them, Platt scaling (?) based approaches are more popular due to their low complexity and efficiency. This includes Temperature Scaling (TS) and its extensions such as attended TS (?), Dirichlet calibration (?), Bin-wise TS (BTS) (?) and Local TS (LTS) (?).
4.1.1 Temperature Scaling
Temperature scaling (TS) is a single-parameter extension of Platt scaling (?) and the most recent addition to the offering of post-hoc methods. It uses a temperature parameter $\tau$ to calibrate the softmax probability:
$p_{i}=\frac{\exp(g_{i}/\tau)}{\sum_{j=1}^{k}\exp(g_{j}/\tau)},~{}~{}i\in[1\dots k].$ (9)
where $\tau>0$ for all classes is used as a scaling factor to soften model predicted probability, it controls the model’s confidence by adjusting the sharpness of distribution so that the model prediction is not too certain (overconfident) or too uncertain (underconfident). The optimal temperature value is obtained by minimizing negative log likelihood loss (NLL) on the validation dataset.:
$\tau^{\ast}=\arg\min_{\tau}\left(-\sum_{i=1}^{N}\log(\textsc{softmax}(g_{i},\tau)) \right)$ (10)
TS simplifies matrix (vector) scaling (?) where class-wise $\tau$ is considered as a single parameter, and offers good calibration while maintaining minimum computational complexity (?, ?).
4.1.2 Temperature Scaling Extensions
The goal of post-hoc calibration on a validation dataset is to learn a calibration map (also known as the canonical calibration function of probability (?)) which transforms uncalibrated probabilities into calibrated ones. Many TS extensions aim to find a proper calibration map.   Kull et al. (?) proposed Dirichlet calibration which assumes probability distributions are parameterized Dirichlet distributions:
$p(x)|y=j\sim Dirichlet(\alpha^{j})$ (11)
where $\alpha^{j}=\{\alpha^{j}_{1},\alpha^{j}_{2},...,\alpha^{j}_{k}\}$ are Dirichlet parameters for $j$-th class. The proposed Dirichlet calibration map family coincides with the Beta calibration family (?). Besides, it provides uniqueness and interpretability as compared to generic canonical parametrization.    (?) suggested that TS has difficulties in finding optimal $\tau$ when the validation set has a limited number of samples. They proposed attended temperature scaling (ATS) to alleviate this issue by increasing the number of samples in validation set. The key idea is to gather samples from each class distribution. Let’s assume $p(y|x)$ is the predicted probability. ATS first divides the validation set into $K$ subsets with $p(y=k|x),k\in[1,K]$, which allows to add more $y\neq k$ samples using the Bayeisan Theorem (?) as the selection criterion:
$p(x,y=k)=\frac{p(y=k|x)}{p(y\neq k|x)}p(x,y\neq k)$ (12)
It indicates that ATS selects the samples with $y\neq k$ which are more probable to belong to $p(x,y=k)$.   Bin-wise TS (BTS) (?) was proposed to extend TS to multiple equal size bins by using the confidence interval-based binning method. Together with data augmentation, BTS showed superior performance as compared to TS. Local TS (LTS) (?) extends TS to multi-label semantic segmentation and makes it adaptive to local image changes. A local scaling factor is learned for each pixel or voxel.
$\mathbb{Q}(x,\tau_{i}(x))=max_{l\in L}\textsc{Softmax}(\frac{g_{i}(x)}{\tau_{i}(x)})^{(l)}$ (13)
where $l$ is the class index, $g_{i}(x)$ is the logit of input at location $x$, and The $\tau_{i}(x)$ is the location-dependent temperature.
Remarks: Although using a single global hyper-parameter, TS remains a popular approach due to its effectiveness and accuracy-preserving (?). Post-hoc calibration usually works well without a huge amount of validation data and is thus data efficient. The calibration procedure is decoupled from training and does not introduce training complexity. On the other hand, post-hoc calibration is less expressive and suffers difficulty in approximating the canonical calibration function when data is not enough (?).
4.2 Regularization Method
Regularization is important to prevent neural network models from overfitting. In this section, we discuss some representative work in this direction that either explicitly or implicitly regularizes modern neural networks to have better calibration.
4.2.1 Explicit Regularization
The typical (explicit) way to add regularization term (or penalty) $L_{reg}$ to standard loss objective (e.g., negative log likelihood):
$\mathcal{L}(\theta)=-\sum\log p(y|x)+\alpha L_{reg}.$ (14)
where $\alpha$ controls the importance of penalty in regularizing weight $\theta$. L2 regularization has been widely used in training modern neural networks and showed its effectiveness in model calibration ](?). Entropy regularization (?):
$\mathcal{L}(\theta)=-\sum\log p(y|x)-\alpha H(p(y|x)).$ (15)
directly penalizes predictive distributions that have low entropy, prevents these peaked distributions, and ensures better model generalization. Recently, (?) analyzed some current calibration methods such as label smoothing (?), focal loss (?), and explicit confidence penalty (?) from a constrained-optimization perspective and pointed out those can be seen as approximations of a liner penalty. The authors proposed to add a margin $m_{a}$ into learning objective:
$\min\mathcal{L}_{CE}+\alpha\sum_{k}\max(0,\max_{j}(g_{j})-j_{k}-m_{a})$ (16)
(?) proposed to directly add the difference between confidence and accuracy (DCA) to standard cross-entropy as a penalty term to improve calibration. (?) further extend this non-differentiable term to a differentiable version using Huber loss, which is used as an explicit regularisation. They also proposed to use a hyper-parameter to control the transition of explicit L1 and L2 regularizers. The Huber loss is then combined with focal loss which provides implicit regularisation effects.
4.2.2 Implicit Regularization: Focal Loss and Its Extensions
Focal loss (?) was originally proposed to alleviate the class imbalance issue in object detection:$\mathcal{L}_{f}=-\sum_{i=1}^{N}(1-p_{i})^{\gamma}\log p_{i}$ where $\gamma$ is a hyperparameter. It has been recently shown that focal loss can be interpreted as a trade-off between minimizing Kullback–Leibler (KL) divergence and maximizing the entropy, depending on $\gamma$ (?):
$\mathcal{L}_{f}\geq\textsc{KL}(q\parallel p)+\mathbb{H}(q)-\gamma\mathbb{H}(p)$ (17)
The first term pushes the model to learn a probability $p$ to have a high value (confident, as close as possible to ground-truth label distribution, which is usually one-hot representation). The second term is constant. The last term regularizes probability to not be too high (overconfident).   Mukhoti et al. (?) empirically observed that $\gamma$ plays an important role in implicitly regularizing entropy and weights. However, finding an appropriate $\gamma$ is challenging for all samples in the datasets. Thus, they proposed sample-dependent scheduled $\gamma$ (FLSD) based on the Lambert-W function (?). They have shown that scheduling $\gamma$ values according to different confidence ranges helps to improve model calibration on both in-domain and out-of-domain (OOD) data.
4.2.3 Implicit Regularization: Differentiable Calibration Proxy
Recall that the aforementioned methods use a penalty term (either explicitly or implicitly) to improve model calibration on dataset $D$. There is a rising direction that directly optimizes objective function by using calibration errors (CE) as differentiable proxy  (?, ?) to standard loss:
$\arg\min_{\theta}L_{standard}(D,\theta)+L_{calibration}(D,\theta)$ (18)
The focus of this line of work is to find differentiable approximations to calibration errors.   Kumar et al.  (?) proposed a kernel-based approach to explicitly calibrate models in training phrase called Maximum Mean Calibration Error (MMCE), which is differentiable and can be optimized using batch stochastic gradient algorithms. They cast the calibration error to be differentiable by defining an integral probability measure over functions from a reproducing kernel Hilbert space (RKHS) $\mathcal{H}$ induced by a universal kernel $k(\cdot,\cdot)$ and cannonical feature map $\phi:[0,1]\rightarrow\mathcal{H}$:
$\textsc{MMCE}(P(r,c))=\left\|E_{(r,c)\sim P}[(c-r)\phi(r)]\right\|_{\mathcal{H}}$ (19)
where $r,c$ represent confidence and correctness scores, respectively, and $P(r,c)$ denotes the distribution over $r,c$ of the predicted probability $P(y|x)$.   An approach that combines meta-learning and a differentiable calibration proxy was proposed by (?). The authors developed a differentiable ECE(DECE) and used it as learning objective for a meta network. The meta network takes representations from original backbone network and outputs a unit-wise L2 weight decay coefficient $\omega$ for backbone network. The DECE is optimized against calibration metrics with validation set but attached to standard cross-entropy (CE) loss.
$\omega^{*}$ $=\arg\min_{\omega}\mathcal{L}_{DECE}^{val}(f_{c}^{*}\circ f_{\theta}^{*}(\omega))$ (20)
$f_{c}^{*},f_{\theta}^{*}(\omega)$ $=\arg\min_{f_{c},f_{\theta}(\omega)}(\mathcal{L}_{CE}^{train}(f_{c}\circ f_{\theta}(\omega)+\omega\left\|f_{c}\right\|^{2})$ (21)
where $f_{c}$ is classification layer and $f_{\theta}$ is the feature extractor.
Mukhoti and Gal (?) proposed patch accuracy versus patch uncertainty (PAvPU) for evaluate uncertain estimates for each image on segmentation task; Motivated by this idea,  (?) extended it to a differentiable proxy, accuracy versus uncertainty calibration (AvUC), which approximations to the accuracy versus uncertainty (AvU):
$AvU=\frac{n_{AC}+n_{IU}}{n_{AC}+n_{AU}+n_{IC}+n_{IU}}$ (22)
The $n_{AC}$, $n_{AU}$, $n_{IC}$, $n_{IU}$ represent the number of samples in the categories of accurate and certain (AC), accurate and uncertain (AU) and certain (IC), inaccurate an uncertain (IU). The AvUC loss is 0 only when all the accurate predictions are certain and inaccurate predictions are uncertain:
$\mathcal{L}_{AvUC}=-\log(\frac{n_{AC}+n_{IU}}{n_{AC}+n_{AU}+n_{IC}+n_{IU}})$ (23)
where $n_{AC}$, $n_{AU}$, $n_{IC}$, $n_{IU}$ are approximated by individual proxy functions. This loss can also be added as regularizer to standard learning objective.
Similarly (?) presented MDCA (Multi-class Difference in Confidence and Accuracy (MDCA)) for improving model calibration. It is a differentiable calibration loss function and independent of the binning mechanism. The idea was inspired by Static Calibration Error (SCE) (?). The authors showed the the combination with focal loss lead the least calibration error.
Remarks: Regularization methods, as compared to post-hoc methods, can directly output a well-calibrated model without additional steps. The increased complexity is different for different methods; for instance, focal loss can be seen as an implicit regularization and does not introduce observable additional computational complexity. Kernel-based and meta-network regularization add additional computations depending on the designed kernel methods and meta-networks.
4.2.4 Implicit Regularization: Data Augmentation
This line of work is highly relevant to regularization methods, instead of directly adding penalty terms to optimization objectives. Those studies try to augment data or add noise to training samples to mitigate model miscalibration. Label smoothing (?) and mixup (?) are popular approaches in this line.   Label smoothing (?) soften hard labels with an introduced smoothing parameter $\alpha$ in the standard loss function (e.g., cross-entropy):
$\mathcal{L}_{c}=-\sum_{k=1}^{K}y_{k}^{s}\log p_{i},~{}~{}y_{k}^{s}=y_{k}(1-\alpha)+\alpha/K$ (24)
where $y_{k}$ is the soft label for $k$-th category. It is shown that LS encourages the differences between the logits of the correct class and the logits of the incorrect class to be a constant depending on $\alpha$. The confidence penalty can be recovered by assuming the prior label distribution is a uniform distribution $u$ and reversing the direction of the KL divergence.
$\mathcal{L}(\theta)=-\sum\log p(y|x)-\textsc{KL}(u\parallel p(y|x)).$ (25)
Mixup training (?) is another work in this line of exploration. It studies the effectiveness of mixup (?) with respect to model calibration (?). Mixup generates synthetic samples during training by convexly combining random pairs of inputs and labels as well. To mix up two random samples $(x_{i},y_{i})$ and $(x_{j},y_{j})$, the following rules are used:
$\bar{x}$ $=\alpha x_{i}+(1-\alpha)x_{j}$ (26)
$\bar{y}$ $=\alpha y_{i}+(1-\alpha)y_{j}$ (27)
where $(\bar{x}_{i},\bar{y}_{i})$ is the virtual feature-target of original pairs. The authors observed that mixup-trained models are better calibrated and less prone to overconfidence in prediction on out-of-distribution and noise data. It is pointed out that mixing features alone does not bring calibration benefits; label smoothing can significantly improve calibration when used together with mixing features.
Remarks: This line of work combats overfitting by data augmentation in hidden space. This improves not only model generalization but also calibration. Those methods don’t significantly increase network complexity but usually require more training time due to more generated or synthesized training samples.
4.3 Uncertainty Estimation
This line of work aims to alleviate model miscalibration by injecting randomness.The popular methods are (1) Bayesian neural networks (?, ?), (2) ensembles (?), (3) Monte Carlo(MC) dropout (?) and (4) Gumbel-softmax (?) based approaches (?, ?). The former three sub-categories have been discussed in recent surveys (?, ?). Figure 4 presents the difference of those methods.
Figure 4: The methods of uncertainty estimation (?). (a) Bayesian neural network; (b) MC dropout; (c) Ensembles; (d) Gumbel-Softmax trick.
4.3.1 Bayesian Neural Networks
Given a learning objective is to minimize negative log likelihood, $\mathcal{L}=-\frac{1}{N}\sum_{i}^{N}\log p(y_{i}|x_{i},w)$. The probability distribution is obtained by Softmax function as:
$p(y_{i}=m|x_{i},w)=\frac{\exp(f_{m}(x_{i},w))}{\sum_{k\in M}\exp(f_{k}(x_{i},w)}.$ (28)
In the inference phase, given a test sample $x^{*}$, the predictive probability $y^{*}$ is computed by:
[…]
As posterior $p(w|D)$ is intractable, we perform approximation by minimizing the Kullback-Lieber (KL) distance. This can also be treated as the maximization of ELBO:
[…]
where $\theta$ are the variational parameters. With the re-parametrization trick (?), a differentiable mini-batched Monte Carlo (MC) estimator can be obtained.
The uncertainty estimation can be done by performing $T$ inference runs and averaging predictions: $p(y*|x*)=\frac{1}{T}\sum_{t=1}^{T}p_{w_{t}}(y^{*}|x^{*},w_{t}).$
4.3.2 MC Dropout, Ensembles and Gumbel-Softmax
By following the above-mentioned strategy, MC-dropout (?), ensembles (?) and Gumbel-softmax sampling (?, ?) introduce randomness in different ways. Then the $T$ in equation (30) corresponds to the number of sets of mask vectors from Bernoulli distribution $\{r^{t}\}_{t=1}^{T}$ in MC-dropout, or the number of randomly trained models in Ensembles, which potentially leads to different sets of learned parameters $\omega=\{\omega_{1},...,\omega_{t}\}$, or the number of sets of sampled attention distribution from Gumbel distribution $\{g^{t}\}_{t=1}^{T}$.
Remarks: This line of work requires multiple inference runs to perform approximations. This increases computational overhead significantly as compared to previous methods. On the other hand, besides model calibration, those methods are primarily proposed for uncertainty quantification and estimation. Therefore, network uncertainty can be captured and measured as well.
4.4 Hybrid Calibration Methods
Beside applying each method independently, we can always have calibration compositions by combining two or more methods. One straightforward way to combine non-post-hoc methods with post-hoc methods. For instance, performing Temperature Scaling (TS) after employing the regularization method and implicit calibration (?, ?). (?) observed that the combination of label smoothing and mixup training significantly improved calibration. While there are several possibilities for combining different approaches, we highlight some interesting compositions.
4.4.1 Ensemble Temperature Scaling (ETS)
(?) gave three important definitions related to calibration properties: accuracy-preserving, data-efficient, and expressive. They pointed out that TS is an accuracy-preserving and data-efficient approach but is less expressive. Ensemble Temperature Scaling (ETS) was proposed to improve TS expressivity while maintaining the other two properties:
$T(z;w,\tau)=w_{1}T(z;\tau)+w_{2}z+w_{3}K^{-1}$ (31)
There are three ensemble components: the original TS $T(z;\tau)=(z_{1}^{\tau^{-1}},...,z_{k}^{\tau^{-1}})/\sum_{k=1}^{K}z_{k}^{\tau^{-1}}$, uncalibrated prediction with $\tau=1$ and uniform prediction for each class $z_{k}=K^{-1}$.
4.4.2 Temperature Scaling with MC Dropout
(?) extended TS to dropout variational inference to calibrate model uncertainty. The key idea is to insert $\tau$ before final softmax activation and insert TS with $\tau>0$ before softmax activation in MC integration:$\hat{p}=\frac{1}{N}\sum_{i=1}^{N}\textsc{softmax}(\frac{f_{w_{i}}(x)}{\tau})$ where $N$ forward passes are performed to optimize $\tau$ with respect to NLL on the validation set. Then the entropy of the softmax likelihood is used to represent the uncertainty of all $C$ classes.
$H(p)=-\frac{1}{\log C}\sum p^{c}\log p^{c},H\in[0,1]$ (32)
Remarks: Appropriately combining different calibration types to some degree can further improve calibration, but, it may also combine their disadvantages, for instance, Ensemble Temperature Scaling (ETS) (?) has increased complexity.
4.5 Discussion and Analysis
We have introduced some representative methods for each line of methodology categorization. In this section, we analyze and describe their strengths and limitations. Additionally, we outline the popular benchmark datasets and backbone models widely used in the literature.
4.5.1 Summary of Strengths and Limitations
Table 1 presents the strengths and limitations of each calibration category. In general, different calibration categories have their own advantages. For methods beyond Post-hoc calibration, the limitations mostly revolve around increased model architectural and computational complexity. While an exact comparison in this regard depends on the specific method selected, a rough complexity ranking could be hybrid method, uncertainty estimation, regularization, and post-hoc method, in descending order of complexity.
Calibration Methods Strengthens Limitations
Post-hoc Calibration (1) Simple and effective: this line of work can be easily implemented and integrated into existing models. On the other hand, it shows its effectiveness and accuracy-preserving (?) with highly competitive predictive and calibration performance. (1) Suboptimal performance: Post-hoc calibration techniques may not fully optimize the calibration of the model, potentially leading to suboptimal calibration performance. Additionally, Chidambaram et al., (?) recently suggested that the performance of temperature scaling degrades with the amount of overlap between classes.
Regularization (1) Training-time calibration, as calibration is performed during the training, we can obtain a well calibrated model directly after training, omitting the post-hoc steps.
(2) Alleviating overfitting: regularization methods improves calibration through alleviated overfitting such as L2 regularization and Focal loss (?). Thus this line of methods can also improve generalization performance.
(1) Implementation and Tuning Complexity. Implementing regularization method is less straightforward as compared to post-hoc ones. Also this line of work usually introduce hyper-parameters, for example, smoothing factor in label smoothing (?) and  (?) empirically showed that the importance of $\gamma$ in calibrating deep models. However, selecting the appropriate hyper-parameters can be challenging and may require extensive hyper-parameter tuning.
Uncertainty Estimation (UE) (1) Uncertainty quantification. Bayesian-method based calibration can quantify the uncertainty in the calibrated model (?). (2) Incorporating prior information. UE methods allow incorporation of prior knowledge or beliefs about the parameters being estimated and provide more robust estimates, especially when dealing with limited data (?). (1) Computational complexity. Many uncertainty estimation methods involve computationally intensive procedures, which can be impractical for large datasets or real-time applications, limiting their scalability. For example MC-dropout (?) requires multiple inference runs and perform averaging on predictions.
Hybrid Methods (1) Combines strengths: Hybrid methods can leverage the advantages of multiple approaches, potentially leading to better overall performance by addressing the limitations of individual techniques. Section 4.4 describes some of these methods (1) Increased complexity: Hybrid methods usually introduce additional complexity in model calibration, Integrating multiple techniques into a cohesive framework may pose implementation challenges and increase the risk of errors requiring more extensive validation and testing to ensure robustness and effectiveness.
Table 1: The summarization of strengthens and limitations of each line of calibration methods.
4.5.2 Backbone Models and Datasets
In the literature, the datasets and backbone models (the models to be calibrated) used in calibration methods vary from method to method, posing a challenge for future methods in making comparisons with previous ones. Here, we outline some representative datasets and backbone models that have been utilized in previous works to facilitate future baseline selections.
  • Backbone Models. The widely used backbone models include: LeNet-5 (?), ResNet (?), Fully-connected convolution network (FCN) (?), WideResNet (?), DenseNet (?), ResNext (?), Global-pooling CNN (GP-CNN) (?), LSTM (?).
  • Datasets. The widely used vision datasets are CIFAR10/100 (?), Cambridge-driving Labeled Video Database (CamVid) (?), CoCo (?), LPBA40 (?), Caltech-UCSD Birds (CUB) (?, ?),Stanford Cars (?), SVHN (?), Tiny-ImageNet (?), STL-10 (?), Pascal-Voc-2012 (?), MINST and Fashion-MNIST (?); NLP datasets such as MR (?), TREC (?), Stanford Sentiment Treebank (SST) (?), 20Newsgroups (?), IMDB (?), Penn Treebank (?), WikiText-2 (?)
Table 2 outlines the categorization, datasets and backbone models to be calibrated. As we can see, for backbones, ResNet (?), WideResNet (?) and DenseNet (?) are most popular ones for image classification task. And LSTM (?) and GP-CNN (?) are widely used in NLP tasks. Regarding datasets, mostly all selected methods used CIFAR10/100 (?) for image classification, and 20Newsgroups (?), IMDB (?)
Methods Categorization Measurement Backbone Models Datasets
Dirichlet Calibration (?) Post-hoc ECE,NLL,MCE, CECE, Bier ResNet-110/110SD/152SD, DenseNet-40, WisdeNet-32, LeNet-5 CIFAR10/100, SVHN
ATS (?) Post-hoc ECE,NLL ResNet52, WideResNet-32, DenseNet-40/100, VGG-16 CIFAR10/100, MINST
BTS (?) Post-hoc ECE CNN,ResNet-50/110, WideResNet-28/10, DenseNet-100, VGG-16 CIFAR10/100, CUB, Stanford Cars
LTS (?) Post-hoc ECE,MCE,ACE, SCE FCN, ResNet-101 CoCo, CamVid, LPBA40
FLSD   (?) Regularization ECE,MCE,NLL, ACE, CECE ResNet-50/110, Wide-ResNet-26-10, DenseNet-121 CIFAR10/100, SST, 20Newsgroups
MMCE (?) Regularization ECE, NLL, Brier ResNet-50/110, Wide-ResNet-28-10, Inception-v3, Global Pooling CNN, LSTM, Tree-LSTM CIFAR10/100, 20Newsgroups,IMDB, SST, Birds CUB 200
Meta-Calibration (?) Regularization ECE,NLL,ACE, CECE ResNet-18/50/110, WideResNet-26-10, GP-CNN CIFAR10/100,SVHN, 20Newsgroups
Margin-based LS (?) Regularization ECE, ACE ResNet-50/101,GP-CNN CIFAR10,Tiny-ImageNet,Pascal-Voc-2012,CUB, 20Newsgroups
Mix-Up (?) Regularization ECE, OE ResNet-18/50, ResNext-101, VGG-16 CIFAR10/100,STL-10,ImageNet,Fashion-MNIST,MR, TREC, IMDB
Inverse Focal Loss (?) Regularization ECE ResNet-32, GP-CNN CIFAR10/100, SVHN, 20Newsgroups
SWA-Gaussian (?) Uncertainty NLL WideResNet-28-10, DenseNet-161, Resnet-152,PreResNet-164, VGG-16, LSTM CIFAR10/100, STL-10, ImageNet, Penn Treebank, WikiText-2
Ensemble (?) Uncertainty ECE, NLL, Bier ResNet-18/34, LeNet-5 CIFAR10/100, MNIST
Mix-n-Match (?) Hybird ECE ResNet-110,DenseNet-40,WideResNet-28-10, WideResNet-52-2, LeNet-5, VGG-19, ResNext-101 CIFAR10/100, ImageNet
TS+MC dropout (?) Hybird ECE, UCE ResNet-18/101, DenseNet-169 CIFAR10/100
Table 2: The state-of-the-art calibration methods and their categorization, datasets and backbone models to be calibrated with those methods.
4.6 Calibrating Pre-trained Large Models
Pre-trained large models, including vision, language, or vision-language models, have been increasingly used in many safety-critical and customer-facing applications. The calibration of those large models has been recently studied and revisited (?, ?).
4.6.1 Large Vision Model Calibration
(?) studied recent state-of-the-art vision models that include vision transformer (?) and MLP-mixer (?). They found out that the model size and amount of pre-training in the recent model generation could not fully explain the observed decay of calibration with distribution shift or model size in the prior model generation. They also discussed the correlation between in-distribution and out-of-distribution (OOD) calibration. They pointed out that the models with better in-distribution calibration also gave better calibration on OOD benchmarks. (?) studied the calibration of CLIP (?) as a zero-shot inference model and found that CLIP is miscalibrated in zero-shot settings.They showed effectiveness of learning a temperature on an auxiliary task and applying it to inference regardless of prompt or datasets.
4.6.2 Large Language Models (LLMs)
The LLMs and prompting engineering have become an efficient learning paradigm and can perform numerous natural language tasks without or with only few examples. However, the outcome of this learning paradigm can be unstable and introduces bias with various prompt templates and training examples (?). The introduced biases include majority label bias, recency bias and token bias. To mitigate this bias, (?) proposed contextual calibration procedure to improve the predictive power of GPT-3 on few-shot learning tasks with a context-free input such as ”N/A”. The contextual calibration is performed by using vector scaling (?).
$\hat{p}=p(y[``N/A"],\textsc{Context}),~{}~{}p_{calibrated}=W\hat{p}$ (33)
where $W$ is a diagonal matrix.
Later on  (?) proposed domain-context calibration (DC) to mitigate label biases in in-context learning (ICL). The method tries to use random words sampled from the task corpus to estimate the effects of different label biases:
$\hat{p}=\frac{1}{T}\sum_{t=1}^{T}p(y[\textsc{In-Domain Random text}]_{t},\textsc{Context})$ (34)
$p_{calibrated}=p(y|x,\textsc{Context})-\hat{p}$ (35)
where $W$ is a diagonal matrix. Similarly, (?) investigated the selection bias of LLMs in multi-choice question (MCQ) task and pinpointed that LLMs’ token bias is an intrinsic cause of the selection bias. They proposed label-free, inference-time debasing method PriDe, which demonstrates notable effectiveness and efficiency, interpretability, and cross-domain transferability. (?) extended mixup (?) training to improve model calibration by synthesizing samples based on the Area Under the Margin (AUM) for pre-trained language models.
Prototypical calibration (PROCA)  (?) is one of the latest studies in calibrating LLMs. It showed the importance of decision boundary in few-shot classification settings and suggested learning a better boundary with a prototypical cluster. Concretely, it estimates $K$ category-wise clusters with the Gaussian mixture model (GMM):
$P_{GMM}(\hat{p})=\sum_{k=1}^{K}\alpha_{k}p_{G}(\hat{p}|u_{k},\Sigma_{k})$ (36)
$p_{calibrated}=\frac{1}{T}\sum_{t=1}^{T}P_{GMM}(\hat{p}_{t})$ (37)
where $u_{k}$ and $\Sigma_{k}$ are the mean vector and covariance matrix of the distribution. The parameters $\{\alpha,u,\Sigma\}_{k=1}^{K}$ are estimated by using the Expectation-Maximization (EM) algorithm (?) with a small unlabelled dataset.
Most recently, (?) identified the limitations of previous approaches, for example, contextual and domain-context calibration have limitations in multi-sentence classification when the estimation of the contextual bias is inaccurate, and prototypical calibration is vulnerable to overfitting problem. Batch Calibration was then proposed to estimate the contextual bias for each class.
4.7 Calibration Applications
In this section, we describe the application of calibration to different relevant domains.
4.7.1 Medical Imaging and Analysis
Calibrated deep learning models play a crucial role in medical imaging tasks such as disease diagnosis and medical image analysis (?). These models provide accurate predictions and uncertainty estimates, assisting clinicians in making informed decisions about patient care. For instance, calibrated models improve the reliability of automated diagnoses in radiology, pathology, and other medical imaging domains. By aligning predicted probabilities with observed outcomes, these models enhance the trustworthiness of automated medical decision support systems. Numerous studies have demonstrated the effectiveness of model calibration techniques in improving the accuracy and interpretability of medical imaging models.  (?) presents a network calibration procedure to make be robust to label noise based on observation that the confusion matrix of the noisy labels can be expressed as the matrix product between the confusion matrix of the clean labels and the label noises. (?) identified the issue of discrimination and calibration biases in models trained for automatic detection of malignant dermatological conditions from skin lesions images. They found that calibration metrics may appear to show that model outputs do not fairly represent uncertainty.
4.7.2 Financial Decision Making
In the finance industry, calibration can provide more reliable access to financial risk assessment such as fraud detection (?, ?), and financial forecasting (?). These models produce accurate predictions of asset prices, market trends, and financial risks, assisting investors and financial institutions in decision-making and risk management. For example, calibrated models help financial institutions identify fraudulent transactions more effectively by providing reliable estimates of transaction risk. Additionally, calibrated models improve the accuracy of financial forecasting models, enabling investors to make informed decisions about portfolio allocation and asset management. Several studies have demonstrated the effectiveness of model calibration techniques in improving the reliability and accuracy of financial forecasting models.
4.7.3 Autonomous Driving
In the domain of autonomous vehicles, calibrating model predictions is essential for safe and reliable navigation in complex traffic scenarios. These models enable accurate perception, prediction, and decision-making, ensuring that self-driving vehicles can operate safely in dynamic environments. For example,  (?) applied MC-dropout (?) to quantify uncertainty of object detection and performed calibration based on the uncertainty analysis.
5 Conclusion and Future Work
We have reviewed the state-of-the-art calibration methods, described with the motivations, causes, measurement metrics, and categorizations. Then we discussed the details and principles of recent methods as well as their individual advantages and disadvantages. Despite recent advances in calibrating deep models, there are still some challenges and under-explored aspects and needs further exploration.
5.1 Mitigating Calibration Bias
Accurately and reliably measuring calibration is still challenging due to the introduced biases from the binning mechanism and the finite sample numbers (?). For the former challenge, it mainly suffers from sensitivity and data inefficiency issues. The sensitivity to the binning scheme is presented in 5. We can see that for a given model, increasing bin numbers gives higher ECE and MCE scores. A KDE-based ECE Estimator (?) was proposed to replace histograms with non-parametric density estimators, which are continuous and more data-efficient. Measuring the bias is then important for having a correct calibration evaluation. Roelofs et al. (?) proposed Bias-by-Construction (BBC) to model the bias in bin-based ECE as a function of the number of samples and bins. It confirms the existence of non-negligible statistical biases. To follow this line of work, future efforts will include developing an unbiased calibration estimator, exploring the trade-off between calibration bias and variance as mentioned in  (?).
Figure 5: The reliability diagrams for a model trained on CIFAR100 with different bin numbers (left to right: 20, 50, 100). The diagonal dash presents perfect calibration, the red bar presents the gap to perfect calibration on each bin.The calibration error is sensitive to increasing bin numbers.
5.2 Calibrating Generative Models
Recent efforts in model calibration have predominantly centered around classification and regression tasks, leaving a noticeable gap in discussions regarding calibration for sequence generation. While advancements in calibration techniques have greatly benefited these tasks, the applicability and efficacy of such techniques in sequence generation remain under-explored. Building upon the seminal work of  (?), which highlighted the inadequately calibrated token-level probabilities in neural machine translation tasks, there is a pressing need to extend these insights to a broader range of generative models. The findings of (?, ?) revealed a crucial connection between token-level probability miscalibration and the counter-intuitive drop in BLEU scores with increased beam size, as noted by (?). Furthermore, they demonstrated that improved sequence-level calibration can be achieved through the re-calibration of token-level probabilities. This underscores the importance of addressing token-level probability miscalibration not only for enhancing performance metrics but also for improving the overall reliability and coherence of generated sequences.
Expanding upon these observations, recent studies such as that by (?) have confirmed the presence of token-level probability miscalibration in Large Language Models (LLMs) utilized for few-shot learning tasks. The identified culprit, the softmax bottleneck as elucidated by (?), particularly on a large vocabulary, further exacerbates this issue. In the context of sequence generation, early token probability miscalibration can have cascading effects, magnifying errors throughout the entire sequence.
Moving forward, addressing the challenge of effectively calibrating token-level probabilities in various settings represents a promising avenue for research, especially in the era of LLMs. Extending the findings from classification and regression tasks to sequence generation domains requires innovative approaches tailored to the unique characteristics and challenges posed by generative models. Insights gained from recent advancements such as generative calibration (?), sequence likelihood calibration (?) and its extension (?).